# AgriField3D: A Curated 3D Point Cloud and Procedural Model Dataset of Field-Grown Maize from a Diversity Panel

### v1

@@ Abstract
The application of artificial intelligence (AI) in three-dimensional (3D) agricultural research, particularly for maize, has been limited by the scarcity of large-scale, diverse datasets. While 2D image datasets are abundant, they fail to capture essential structural details such as leaf architecture, plant volume, and spatial arrangements that 3D data provide. To address this limitation, we present AgriField3D (website), a curated dataset of 3D point clouds of field-grown maize plants from a diverse genetic panel, designed to be AI-ready for advancing agricultural research. Our dataset comprises over 1,000 high-quality point clouds collected using a Terrestrial Laser Scanner, complemented by procedural models that provide structured, parametric representations of maize plants. These procedural models, generated using Non-Uniform Rational B-Splines (NURBS) and optimized via a two-step process combining Particle Swarm Optimization (PSO) and differentiable programming, enable precise, scalable reconstructions of leaf surfaces and plant architectures. To enhance usability, we performed graph-based segmentation to isolate individual leaves and stalks, ensuring consistent labeling across all samples. We also conducted rigorous manual quality control on all datasets, correcting errors in segmentation, ensuring accurate leaf ordering, and validating metadata annotations. The dataset further includes metadata detailing plant morphology and quality, alongside multi-resolution subsampled versions (100k, 50k, 10k points) optimized for various computational needs. By integrating point cloud data of field grown plants with high-fidelity procedural models and ensuring meticulous manual validation, AgriField3D provides a comprehensive foundation for AI-driven phenotyping, plant structural analysis, and 3D applications in agricultural research.
@@ 1 Introduction
In the past few years, there has been increasing interest in creating 3D plant datasets to address phenotyping needs. However, creating comprehensive real-world 3D datasets presents significant challenges, especially in agriculture. The process requires substantial resources, including field and greenhouse space, expert agricultural knowledge, and extended periods for plant growth cycles, making it both expensive and complex. To overcome these limitations, researchers have explored synthetic alternatives that can generate large-scale datasets efficiently. For example, Zarei et al. [22] utilized procedural modeling to generate 64,000 annotated sorghum plants, while Anagnostopoulou et al. [23] created 15,000 annotated mushroom images using generative models. While this virtual generation method offers a promising alternative for creating extensive datasets, the synthetic data may not fully capture the nuances of real plants, especially under field-grown conditions, potentially limiting its practical applications.
@@
AgriField3D (Ours)  1,045 (Maize) TLS, Procedural models Yes Semantic & Instance. Includes 520 annotated point clouds with segmentation for leaves and stalks. Provides multiple subsampled datasets (100k, 50k, 10k) tailored for different downstream applications. High-resolution AI-ready data ideal for phenotypic analysis and ML applications.
Table 1 provides a comparison of existing publicly available 3D plant datasets, highlighting their key characteristics and limitations. These datasets vary widely in size, scanning technology, and annotation detail. The smallest dataset, ROSE-X, includes 11 rose plant models, while Crops3D is the largest with 1,230 3D plant point clouds across eight crop species. Among single-crop datasets, the Single Maize Point Cloud dataset, with 428 maize samples, represents the largest publicly available collection dedicated to a single species.
@@
Pheno4D [24] introduces a dataset featuring 7 maize plants and 7 tomato plants. These plants were scanned daily using a 3D laser scanning system over two weeks for maize and three weeks for tomato plants, resulting in 84 point clouds for maize (with 49 labeled) and 140 point clouds for tomato (with 77 labeled). The dataset does not include color. Labels applied to the dataset include semantic and instance labels. ROSE-X [25] provides a dataset of 11 complete 3D models of real rosebush plants. These models were acquired using X-ray computed tomography (CT), offering voxel-level annotations for the entire plant shoot, with organ-level labels for stem, leaf, and flower. While the organ labels are color-coded in the point cloud representations, the dataset does not specifically include true color information. A dataset of 428 maize samples was introduced by [26], where point clouds were collected using a 3D laser scanner. These samples spanned five maize varieties (Xian Yu 335, LD145, LD502, LD586, and LD1281) and were annotated using the Label3DMaize toolkit for semantic and instance segmentation. Downsampled point clouds were processed to a uniform size of 20,480 points per sample, ensuring efficient use in training. A total of 709 plant shoot architectures from four species (Arabidopsis, Tomato, Tobacco, and Sorghum) are featured in Plant3D [27, 28, 29]. These architectures have been scanned across 20–30 developmental time points and under a variety of situations, including ambient light, high heat, high light, vegetative shade, and drought. Soybean-MVS [30] introduces a dataset consisting of 102 soybean 3D plant models reconstructed using multiple-view stereo (MVS) technology. The dataset spans the entire soybean growth period and includes five varieties. Color data was included in the point clouds (xyzRGB format). The dataset features semantic labels for plant organs, including leaves, main stems, and stems. A dataset of 34 actual plant point clouds from three species—Capsicum annuum (pepper), Rosa kordana (rose), and Ribes rubrum (ribes)—is presented by PLANesT-3D [31]. These plants were reconstructed using Structure from Motion (SfM) and MVS techniques, offering annotated 3D color point clouds. Points were labeled as either ”leaf” or ”stem,” and individual leaflets within the plants were uniquely labeled. Crops3D [32] provides a comprehensive dataset comprising 1,230 colored 3D crop point cloud samples from eight crop species including cabbage, cotton, maize, potato, rapeseed, rice, tomato, and wheat. Point clouds were acquired using a combination of TLS, Structured Light Scanning, and SfM-MVS techniques. Points were categorized into specific plant organs such as soil, stem, leaf, tassel, fruit, pot, and others, depending on the crop. The dataset includes instance annotations for individual plants within plots.
@@
While these datasets have advanced plant-specific ML algorithms and segmentation methods, they are often limited by challenges such as low-resolution models, incomplete annotations, small dataset sizes, and restricted diversity of plant conditions. For maize, existing datasets are small in scale, lack comprehensive annotations, or fail to incorporate field-grown plants representing real-world phenotypic variability. These gaps hinder the development of robust AI-driven phenotyping solutions, particularly for traits like organ-level segmentation and trait analysis.
@@
To overcome these limitations, we introduce AgriField3D, a curated, large-scale dataset with 1,045 high-resolution 3D point clouds of field-grown maize plants captured using TLS and additionally represented using procedural modeling techniques. This dataset was collected from the SAM diversity panel. It includes 520 annotated point clouds with semantic and instance segmentation for leaves and stalks. Additionally, AgriField3D incorporates procedurally generated maize models, which were created for each of the 520 segmented maize plants as described in Section 3.1.5. This procedural modeling approach allows for low dimensional, surface based representation of realistic maize canopies by leveraging actual plant data. The procedural models provide advantages for AI approaches including data augmentation, and synthetic-to-real transfer learning. To further support AI-driven research, AgriField3D provides multiple subsampled versions (100k, 50k, and 10k points per cloud) optimized for different computational needs.
The main contributions of AgriField3D are:
  • 1,045 quality-controlled 3D point clouds of diverse field-grown maize plants.
  • Organ-level annotations with detailed semantic and instance segmentation, enabling precise plant structure analysis.
  • AI-compatible formats, ensuring seamless integration with advanced ML models.
  • Color information, enhancing data interpretability.
  • Procedurally generated maize leaf models, providing a NURBS surface-based representation that is beneficial for downstream tasks.
  • Multiple subsampled versions (100k, 50k, and 10k points per cloud), supporting a range of computational requirements.
Our dataset is among the largest and most detailed resources for maize phenotyping, providing a robust foundation for training and validating AI-based phenotyping algorithms.
@@
In the following sections, we detail the data processing methodologies, dataset annotation, metadata and quality control, and potential applications and highlight its significance for advancing agricultural research. By providing this dataset, we aim to catalyze the development of innovative AI-driven phenotyping solutions, paving the way for new insights into maize biology and crop improvement.
@@ 2.1 Dataset access
The AgriField3D dataset [^1] https://baskargroup.github.io/AgriField3D/ is publicly available through the Hugging Face Datasets platform [^2] https://huggingface.co/datasets/BGLab/AgriField3D, ensuring accessibility and ease of integration into machine learning pipelines. The dataset includes high-resolution 3D point clouds and segmented plant models, organized for efficient navigation and usability. Users can download the complete dataset or select specific components tailored to their research needs. Key features include:
  • Point Clouds: Point cloud data of over 1,000 maize plants at multiple subsampled resolutions ($10,000$, $50,000$, and $100,000$ points) to accommodate various computational requirements.
@@
  • Documentation: A detailed user guide is included in the repository, offering step-by-step instructions for integrating the dataset with tools such as PyTorch, TensorFlow, and Scikit-learn, enabling streamlined machine learning applications.
@@
Figure 1: Nine examples of maize plant point clouds from the dataset, showcasing the original data. These images highlight the diverse morphologies of maize plants captured using terrestrial laser scanning.
@@ 2.2 Dataset Composition
The dataset comprises over 1000 high-resolution point clouds, each capturing detailed 3D information of individual maize plants. The dataset represents a diverse range of genetic backgrounds and provides the following key components:
@@
  • Around 520 segmented plants into stalks and leaves with specific color coding, ordered from bottom to top leaves.
  • Metadata detailing plant quality, leaf count, presence of tassels and ears, and other critical attributes.
  • Multiple resolutions to support diverse AI applications.
Figure 1 showcases nine example maize plants from the dataset, highlighting the original data without segmentation and the diversity in plant morphology.
@@ 2.3 Data Collection
Field-grown maize plants harvested at about the time of anthesis were selected for data collection from the maize SAM (shoot apical meristem) diversity panel [33] at Iowa State University’s Curtis Farm. These lines were planted in 4-row plots at the Iowa State University Woodruff Farm in Ames, Iowa, in the year 2021. Each plot measured 3.04 meters in row length, with plants spaced approximately 15.24 cm apart within rows and 76.2 cm apart between rows, resulting in a planting density of around 84,000 plants per hectare. All plants were collected from the field at the time of anthesis and immediately brought to indoor environment for imaging and 3D scanning. Each plant was positioned upright using specialized holders to maintain a natural posture, with barcodes scanned to link each plant to its variety tag. Corresponding tag names were printed on white sheets and placed adjacent to the plants for precise labeling and consistent organization.
The 3D data were captured using a Faro Focus S350 Scanner, depicted in Figure 2(a), which is equipped with an angular resolution of 0.011 degrees, enabling precise point spacing of approximately 1.5 mm at a 10-meter range. The scanner has the capacity to acquire up to 700 million points at a rate of 1 million points per second, ensuring high-resolution outputs suitable for detailed phenotypic analysis. Plants were arranged in batches of eight around the scanner in a circular configuration, as shown in Figure 3, allowing simultaneous capture of multiple samples. This setup minimized overlaps between plants and ensured most surfaces were within the scanner’s line of sight. The scanner height and settings were dynamically adjusted based on plant height and morphology, ensuring that critical structural features, including upper leaves, stalks, tassels, and ears, were precisely captured. Figure 2(b) provides a visual representation of the scanning setup, showcasing the Faro scanner in operation while capturing maize plants arranged in a circular configuration. The controlled indoor environment eliminated external variables such as wind and inconsistent lighting, ensuring uniformity and accuracy across the dataset. Scans were conducted batch by batch, until the full dataset was acquired.
Figure 2: (a) The Faro Focus S350 Scanner is used for collecting 3D data of maize plants. This advanced scanner ensures precise and high-resolution point cloud generation, which is critical for phenotypic analysis. (b) Visualization of the scanning process using the Faro scanner. Maize plants are arranged in a circular configuration, allowing simultaneous high-resolution capture of multiple samples in a controlled indoor environment.
Figure 3: Point cloud visualization of maize plants. (a) The circular arrangement of eight maize plants mounted on a specialized platform, each labeled with unique identifiers. (b) Detailed view of a single maize plant showing characteristic leaf architecture and stem structure.

### v3

# MaizeField3D: A Curated 3D Point Cloud and Procedural Model Dataset of Field-Grown Maize from a Diversity Panel

@@ Abstract
The development of artificial intelligence (AI) and machine learning (ML) based tools for 3D phenotyping, especially for maize, has been limited due to the lack of large and diverse 3D datasets. 2D image datasets fail to capture essential structural details such as leaf architecture, plant volume, and spatial arrangements that 3D data provide. To address this limitation, we present MaizeField3D (website), a curated dataset of 3D point clouds of field-grown maize plants from a diverse genetic panel, designed to be AI-ready for advancing agricultural research. Our dataset includes 1,045 high-quality point clouds of field-grown maize collected using a terrestrial laser scanner (TLS). Point clouds of 520 plants from this dataset were segmented and annotated using a graph-based segmentation method to isolate individual leaves and stalks, ensuring consistent labeling across all samples. This labeled data was then used for fitting procedural models that provide a structured parametric representation of the maize plants. The leaves of the maize plants in the procedural models are represented using Non-Uniform Rational B-Spline (NURBS) surfaces that were generated using a two-step optimization process combining gradient-free and gradient-based methods. We conducted rigorous manual quality control on all datasets, correcting errors in segmentation, ensuring accurate leaf ordering, and validating metadata annotations. The dataset also includes metadata detailing plant morphology and quality, alongside multi-resolution subsampled point cloud data (100k, 50k, 10k points), which can be readily used for different downstream computational tasks. MaizeField3D will serve as a comprehensive foundational dataset for AI-driven phenotyping, plant structural analysis, and 3D applications in agricultural research.
@@ 1 Introduction
In the past few years, there has been increasing interest in creating 3D plant datasets to address phenotyping needs. However, creating comprehensive real-world 3D datasets presents significant challenges, especially in agriculture. The process requires substantial resources, including field and greenhouse space, expert agricultural knowledge, and extended periods for plant growth cycles, making it both expensive and complex. To overcome these limitations, researchers have explored synthetic alternatives that can generate large-scale datasets efficiently.
One approach involves generating simulation-ready 3D plant datasets and systematically constructing CAD models to populate virtual fields. Using tools like SolidWorks and Blender, these models can be embedded into configurable simulation environments (e.g., Gazebo), facilitating visual navigation and field mapping experiments for autonomous phenotyping platforms. Such environments allow flexible modeling of real-world field conditions with configurable crop and weed types, layouts, and occlusions using mesh libraries and randomized pose distributions  [22]. Although mesh or CAD-based synthetic datasets are scalable and customizable, they often lack the realism and variability found in actual agricultural scenes. To improve the generalizability of models trained on synthetic data, researchers have increasingly adopted domain adaptation strategies such as Real2Sim  [23], where real-world measurements guide synthetic model generation. For instance,  [24] proposed a closed-loop Real2Sim framework that uses apple tree point clouds to generate structurally accurate 3D synthetic trees, enabling improved trait extraction under zero-shot conditions without additional real-world data. A recent and promising direction leverages diffusion-based generative models to create visually realistic plant scenes from textual or structured inputs. For instance, Anagnostopoulou et al. [25] developed a diffusion-based framework to synthesize 15,000 photorealistic images of mushrooms, capturing a wide range of natural variability in lighting, background, and shape.
Complementing these efforts, procedural generation techniques offer a flexible means to synthesize complex, variable plant geometries at scale. These methods use rule-based or stochastic modeling approaches to generate diverse plant morphologies and environments without the need for dense real-world scans. For example, Zarei et al. [26] utilized procedural modeling to generate 64,000 annotated sorghum plants. More recently, Liu et al. [27] introduced a neural hierarchical decomposition framework that reconstructs biologically plausible 3D plant structures from single 2D images. By learning to infer a tree-structured graph of organs and their geometric properties, the method enables high-fidelity, simulation-ready plant models from minimal inputs, offering a scalable approach for dataset augmentation and trait analysis.
Our work follows a similar direction. Rather than generating plants entirely from scratch, we reconstruct plant geometry using procedural NURBS surfaces derived from scans of the real field-grown plants. This allows us to maintain biological realism while also enabling scalable generation of synthetic data for tasks like canopy-level light simulation, phenotype extraction, and ML applications. By rooting the synthetic models in real-world structure, we aim to offer a middle ground between pure simulation and costly manual annotation.
@@
MaizeField3D (Ours) 1,045 (Maize) TLS, Procedural models Yes Semantic & Instance. Includes 520 annotated point clouds with segmentation for leaves and stalks. Provides multiple subsampled datasets (100k, 50k, 10k) tailored for different downstream applications. High-resolution AI-ready data ideal for phenotypic analysis and ML applications.
@@
Table 1 provides a comparison of existing publicly available 3D plant datasets, highlighting their key characteristics and limitations. These datasets vary widely in size, scanning technology, and annotation detail. The smallest dataset, ROSE-X, includes 11 rose plant models, while Crops3D is the largest, with 1,230 3D plant point clouds across eight crop species. Among single-crop datasets, the Single Maize Point Cloud dataset, with 428 maize samples, represents the largest publicly available collection dedicated to a single species. Pheno4D [28] introduces a dataset featuring 7 maize plants and 7 tomato plants. These plants were scanned daily using a 3D laser scanning system over two weeks for maize and three weeks for tomato plants, resulting in 84 point clouds for maize (with 49 labeled) and 140 point clouds for tomato (with 77 labeled). The dataset does not include color. Labels applied to the dataset include semantic and instance labels. ROSE-X [29] provides a dataset of 11 complete 3D models of real rosebush plants. These models were acquired using X-ray computed tomography (CT), offering voxel-level annotations for the entire plant shoot, with organ-level labels for stem, leaf, and flower. While the organ labels are color-coded in the point cloud representations, the dataset does not specifically include true color information. A dataset of 428 maize samples was introduced by [30], where point clouds were collected using a 3D laser scanner. These samples spanned five maize varieties (Xian Yu 335, LD145, LD502, LD586, and LD1281) and were annotated using the Label3DMaize toolkit for semantic and instance segmentation. Downsampled point clouds were processed to a uniform size of 20,480 points per sample, ensuring efficient use in training. A total of 709 plant shoot architectures from four species (Arabidopsis, Tomato, Tobacco, and Sorghum) are featured in Plant3D [31, 32, 33]. These architectures have been scanned across 20–30 developmental time points and under a variety of situations, including ambient light, high heat, high light, vegetative shade, and drought. Soybean-MVS [34] introduces a dataset consisting of 102 soybean 3D plant models reconstructed using multiple-view stereo (MVS) technology. The dataset spans the entire soybean growth period and includes five varieties. Color data was included in the point clouds (xyzRGB format). The dataset features semantic labels for plant organs, including leaves, main stems, and stems. A dataset of 34 actual plant point clouds from three species—Capsicum annuum (pepper), Rosa kordana (rose), and Ribes rubrum (ribes)—is presented by PLANesT-3D [35]. These plants were reconstructed using Structure from Motion (SfM) and MVS techniques, offering annotated 3D color point clouds. Points were labeled as either ”leaf” or ”stem”, and individual leaflets within the plants were uniquely labeled. Crops3D [36] provides a comprehensive dataset comprising 1,230 colored 3D crop point cloud samples from eight crop species, including cabbage, cotton, maize, potato, rapeseed, rice, tomato, and wheat. Point clouds were acquired using a combination of TLS, Structured Light Scanning, and SfM-MVS techniques. Points were categorized into specific plant organs such as soil, stem, leaf, tassel, fruit, pot, and others, depending on the crop. The dataset includes instance annotations for individual plants within plots.
@@
While these datasets have advanced plant-specific ML algorithms and segmentation methods, they are often limited by challenges such as low-resolution models, incomplete annotations, small dataset sizes, and restricted diversity of plant conditions. For maize, existing datasets often lack comprehensive annotations or fail to incorporate field-grown plants representing real-world phenotypic variability. These gaps hinder the development of robust phenotyping solutions, particularly for traits like organ-level segmentation and trait analysis.
@@
To overcome these limitations, we introduce MaizeField3D, a curated, large-scale dataset with 1,045 high-resolution 3D point clouds of field-grown maize plants captured using TLS and additionally represented using procedural modeling techniques. This dataset was collected from the SAM diversity panel. It includes 520 annotated point clouds with semantic and instance segmentation for leaves and stalks. Each segmented point cloud includes RGB color labels that encode leaf identity based on vertical position, facilitating organ-level analysis and visualization. Additionally, MaizeField3D incorporates procedurally generated maize models, which were created for each of the 520 segmented maize plants as described in Section 3.1.5. This procedural modeling approach allows for low dimensional, surface based representation of realistic maize canopies by leveraging actual plant data. The procedural models provide advantages for ML approaches, including data augmentation and synthetic-to-real transfer learning. To further support data-driven research, MaizeField3D provides multiple subsampled versions (100k, 50k, and 10k points per cloud) optimized for different computational needs.
The main contributions of MaizeField3D are:
  • A curated dataset of 1,045 high-quality 3D point clouds representing diverse, field-grown maize plants, with 520 of them annotated at the organ level with detailed semantic and instance segmentation for leaves and stalks.
  • Procedurally generated NURBS surface models of 520 maize plants’ leaves, offering a compact, structured, and biologically meaningful representation for downstream applications such as functional-structural analysis and detailed (bio)physics simulation.
  • RGB color labels encoding leaf identity based on vertical position, supporting consistent organ-level visualization and analysis.
  • Standardized data formats and multiple resolution levels (100k, 50k, and 10k points per cloud), enabling flexible integration into machine learning pipelines with varying computational requirements.
  • A demonstrated use case involving canopy-scale radiation interception simulation using HELIOS [37], where virtual field plots constructed from our procedural models are used to compute absorbed photosynthetically active radiation (PAR) under real solar conditions.
@@
In the following sections, we explain the methods for data processing, dataset annotation, metadata creation, and quality control. By providing this dataset, we aim to catalyze the development of innovative phenotyping solutions, paving the way for new insights into maize biology and crop improvement.
@@ 2.1 Dataset Access
The MaizeField3D dataset [^1] https://baskargroup.github.io/MaizeField3D/ is publicly available through the Hugging Face Datasets platform [^2] https://huggingface.co/datasets/BGLab/MaizeField3D, ensuring accessibility and ease of integration into ML pipelines. The dataset includes high-resolution 3D point clouds and segmented plant models, organized for efficient navigation and usability. Users can download the complete dataset or select specific components tailored to their research needs. Key features include:
  • Point Clouds: Point cloud data of 1,045 maize plants at multiple subsampled resolutions ($100,000$, $50,000$, and $10,000$ points) to accommodate various computational requirements.
@@
  • Documentation: A detailed user guide is included in the repository, offering step-by-step instructions for integrating the dataset with tools such as PyTorch, TensorFlow, and Scikit-learn, enabling streamlined ML applications.
@@
Figure 2: (a) Original point clouds of three maize plants, highlighting the diverse morphologies of maize plants captured using terrestrial laser scanning. (b) Corresponding segmented maize plant point clouds, showcasing the color-coded segmentation data.
@@ 2.2 Dataset Composition
The dataset comprises 1,045 high-resolution point clouds, each capturing detailed 3D information of individual maize plants. The dataset represents a diverse range of genetic backgrounds and provides the following key components:
@@
  • 520 segmented plants into stalks and leaves with specific color coding, ordered from bottom to top leaves.
  • Metadata detailing plant quality, presence of tassels and ears, and other critical attributes.
  • Multiple resolutions to support diverse applications.
Figure 2(a) showcases three example maize plants from the dataset, highlighting the original data without segmentation and the diversity in plant morphology. Additional examples are provided in Figure S.1 (Supplementary Materials).
Figure 3: Scanning hardware, setup, and resulting 3D point clouds from the MaizeField3D dataset. (a) The Faro Focus S350 Scanner is used for collecting 3D data of maize plants. This advanced scanner ensures precise and high-resolution point cloud generation, which is critical for phenotypic analysis. (b) Visualization of the scanning process using the Faro scanner. Maize plants are arranged in a circular configuration, allowing simultaneous high-resolution capture of multiple samples in a controlled indoor environment. Point cloud visualization of maize plants are shown in the bottom. (c) The circular arrangement of eight maize plants mounted on a specialized platform, each labeled with unique identifiers. (d) Detailed view of a single maize plant showing characteristic leaf architecture and stem structure.
@@ 2.3 Data Collection
Field-grown maize plants harvested at about the time of anthesis were selected for data collection from the maize SAM (shoot apical meristem) diversity panel [38] at Iowa State University’s Curtis Farm. These lines were planted in 4-row plots at the Iowa State University Woodruff Farm in Ames, Iowa, in the year 2021. Each plot measured 3.04 meters in row length, with plants spaced approximately 15.24 cm apart within rows and 76.2 cm apart between rows, resulting in a planting density of around 84,000 plants per hectare. All plants were collected from the field at the time of anthesis and immediately brought to an indoor environment for imaging and 3D scanning. Each plant was positioned upright using specialized holders to maintain a natural posture, with barcodes scanned to link each plant to its variety tag. Corresponding tag names were printed on white sheets and placed adjacent to the plants for precise labeling and consistent organization.
The 3D data were captured using a Faro Focus S350 Scanner, depicted in Figure 3(a), which is equipped with an angular resolution of 0.011 degrees, enabling precise point spacing of approximately 1.5 mm at a 10-meter range. The scanner has the capacity to acquire up to 700 million points at a rate of 1 million points per second, ensuring high-resolution outputs suitable for detailed phenotypic analysis. Plants were arranged in batches of eight around the scanner in a circular configuration, as shown in Figure 3(c), allowing simultaneous capture of multiple samples. This setup minimized overlaps between plants and ensured most surfaces were within the scanner’s line of sight. While the scanner setup captured most of each plant’s surface area, the backside of some plants was not directly visible due to the single-scan circular arrangement. This resulted in small gaps in the resulting point clouds. However, the configuration was optimized to capture as much structural detail as possible from a single vantage point, balancing coverage, resolution, and throughput. The scanner height and settings were dynamically adjusted based on plant height and morphology, ensuring that critical structural features, including upper leaves, stalks, tassels, and ears, were precisely captured. Figure 3(b) provides a visual representation of the scanning setup, showcasing the Faro scanner in operation while capturing maize plants arranged in a circular configuration. The controlled indoor environment eliminated external variables such as wind and inconsistent lighting, ensuring uniformity and accuracy across the dataset. Scans were conducted batch by batch until the full dataset was acquired.